Star at 8:00pm

Check the trigger in the UTH, the fiber is changed by the spare and the check test is good. The spare fiber is left connected. Original fiber is checked and it is seems working, this fiber is keep it.

Flattening half resolution

Check of the 500x500 int mat acquire and redux yesterday.
Track.num n.modes rmsWFE(nm) intmat notes
20130827_035840 400 96 20130826_150704 not diverge
20130827_040002 500 98 20130826_150704 not diverge
20130827_040226 600 80 20130826_150704 not diverge
20130827_050422 600 79 20130826_150704 not diverge

Noise measurement

Taking a few burst we can see some convection with a different pattern than yesterday, this could be cause by the High bay hatch that is completely open.Measurements of the noise is taken; 1000 measurements

file: 20130827_041100_noise

The telescope crew started to close the hatch.

file: 20130827_042537_noise, 20130827_045655_noise,20130827_050556_noise

Produce file: 20130827_042537_noise

20130827_060000_noise, 20130827_061521_noise,20130827_065600_noise,20130827_100300_noise

Hexalign calibration.

@9:00pm The hexapod started a new calibration. AdOpt >print,hexapod_opt_calib(ave=5)

hexapod position:
x y z Rx Ry Rz  
-0.19 0.93 4.18 -166 -144 0  

Actuator Mapping

IDL> a='/towerdata/meas/adsec_calib/if_functions/newmodes_20130826/MMmatrix.sav'
IDL> make_offload_matrix_proj,a, /ver

then, the index of pixel corresponding to each capsens may be obtained with: q = draw_act_map(reference_mask, act)

FF-MATRIX TUNNING

Command used; print,collect_ff_data(/no)

The script to run the collection of the data didn't work. When the routine rested the shell and the load program, the start up process didn't complete correctly. It is seems that some files; corrupted the startup of the mirror.

@6:00 am the problem is solved and the acquisiton of the FF-Matrix is started

act#279 is showing a spike like residual in the data. to be excluded

IF modes collection for the KL basis

i4d_opt_intmat(600,660,avera=11,FORCE_sc=0.5/0.3)

print, i4d_opt_intmat(i*25,i*25+24, ave=25)
Tracknum modes average amp Notes
20130827_154317 300-324 25    
20130827_154421 325-349 25    
20130827_155106 350-374 25   using the extraforce here produces a RIP!!
20130827_155201 375,399 25    
20130827_155408 400-424 25    
20130827_155504 425-449 25    
20130827_155558 450-474 25    
20130827_155654 475-499 25    
         
         
         
         
         
         
         


-- ArmandoRiccardi - 27 Aug 2013
Topic revision: r5 - 28 Aug 2013, AlfioPuglisi
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback